### v1

# Ramanujan’s function on small primes

Abstract.
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number. For this reason, as an example in a recent article about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit regularities of the following kind: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in a nearly periodic fashion with $n$.
Key words and phrases:
partitions, graphs, cusp forms
2020 Mathematics Subject Classification:
05A17, 05C62, 11F11, 11P83
1 December 2025
1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] . For this reason, as an example in a recent article [Bre25a] about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{1}=2$ and, more generally, $p_{n}$ denotes the $n^{th}$ prime. We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit rough regularities: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in an approximately periodic fashion with $n$. For obvious reasons, we do not expect that these regularities are sustained without modification throughout the sequence $\{\tau(p_{n})\}$ or any of the other sequences we have examined.
2. Lemmas and definitions
To introduce the matrices that we will associate to the members of the sequence $\{\tau(p_{n})\}_{n\geq 1}$, we repeat some definitions and lemmas from our earlier article [Bre25a]. Although we supplied our own proofs of the lemmas (omitted here), they are well known in the theory of symmetric functions [M]. In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
Lemma 2.1.
The equations below are equivalent. (We will refer to both of them as equation (D) in the sequel.) Let $h_{0}=1$ and
$nh_{n}=\sum_{r=1}^{n}j_{r}h_{n-r}$
for $n\geq 1$. With $H(t)=\sum_{n=0}^{\infty}h_{n}t^{n}$ and $J(t)=\sum_{r=1}^{\infty}j_{r}t^{r}$,
$t\frac{d}{dt}H(t)=H(t)J(t).$
With $f_{0}=1$, let $\overline{f}$ denote the sequence $f_{0},f_{1},f_{2},$.... Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
$\left(\begin{matrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-1}&j_{n-2}&\cdots&j_{1}&-n+1\\ j_{n}&j_{n-1}&j_{n-2}&\cdots&j_{1}\\ \end{matrix}\right)$
and
$\left(\begin{matrix}h_{1}&1&0&\cdots&0\\ 2h_{2}&h_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ nh_{n}&h_{n-1}&h_{n-2}&\cdots&h_{1}\\ \end{matrix}\right),$
respectively.
Lemma 2.2.
Let $h_{0}=j_{1}=1$. Equation (D) implies that
(1) $j_{n}=(-1)^{n+1}|H_{n}(\overline{h})|.$
and
(2) $n!h_{n}=|J_{n}(\overline{j})|$
for $n\geq 1$.
We define deformations $J^{(c)}_{n}(\overline{j})$ of the $J_{n}(\overline{j})$ as the matrices
$\left(\begin{matrix}c&-1&0&\cdots&0\\ j_{1}&c&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-2}&j_{n-3}&\cdots&c&-n+1\\ j_{n-1}&j_{n-2}&j_{n-3}&\cdots&c\\ \end{matrix}\right).$
Remark 1.
At present we have no helpful insight into the theory of these deformed matrices except to say that (obviously) $J^{(c)}_{n}(\overline{j})=J_{n}\left(\overline{j}^{(c)}\right)$ where $\overline{j}^{(c)}=\{c,j_{1},j_{2},$...\}.
Now we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999]. Let $A_{n}=$
$\begin{pmatrix}a_{1}&-1&0&\cdots&0\\ a_{2}&a_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n-1}&a_{n-2}&\cdots&a_{1}&-(n{-}1)\\ a_{n}&a_{n-1}&\cdots&a_{2}&a_{1}\end{pmatrix}^{T},$
where the $T$ operator is matrix transpose. Let $B_{n}(x)=|A_{n}-xI_{n}|$, so that $B_{n}(x)=(-1)^{n}\chi_{A_{n}}(x)$. Then
Proposition 1.
(Vein and Dale, Theorem 4.23)
$B_{n}(x)=\sum_{r=0}^{n}\binom{n}{r}|A_{r}|x^{n-r}.$
Theorem 2.3.
 [^1] Here we are merely codifying a method shown to us by J. López-Bonilla. (See Corollary 3.3 in our previous paper [Bre25a].)
With $h$ and $j$ as in the above lemmas,
$\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
Proof.
Determinants are invariant under the transpose, so, setting each $a_{i}$ equal to $j_{i}$ in the lemmas above, the proposition can be restated in the following way: Let $J_{n}(\overline{j})=$
$\begin{pmatrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ j_{n-1}&a_{n-2}&\cdots&j_{1}&-(n{-}1)\\ j_{n}&j_{n-1}&\cdots&j_{2}&j_{1}\end{pmatrix}$.
Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. Then $C_{n}(x)=$
$\sum_{r=0}^{n}\binom{n}{r}|(J_{r}(\overline{j})|(-1)^{n-r}x^{n-r}=$
(by Lemma 2.2)
$\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-x)^{n-r}=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{n-r}x^{n-r}.$
Simplifying, $\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
3. Empirical observations about the sequence $\{\tau(p_{n})\}_{n\geq 1}$
3.1. Characteristic functions of the undeformed matrices representing the function $n\mapsto\tau(p_{n})$
Let $h_{\tau,p}(0)=1$ and $h_{\tau,p}(n)=\tau(p_{n})$ for positive integers $n$. Let $j_{\tau,p}(n)$ and $h_{\tau,p}(n)$ satisfy equation (D) together. Clearly, the values of $j_{\tau,p}(n)$ can be computed from equation (D1). Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau,p}}))=|xI-J_{n}(\overline{j_{\tau,p}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau,p}})$. It is clear that the following statements are equivalent: (a) $\tau(p_{n})=0$, (b) $|J_{n}(\overline{j_{\tau,p}})|=0$, and (c) the polynomial $\chi(J_{n}(\overline{j_{\tau,p}}))$ takes zero to zero.
In the range of our observations, the degree of $\Pi_{n}(x)$ is $n$. Writing $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k},$ [^2] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25].
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau(p_{k})$
(also in the range of our observations.) Equation (3) is immediate after Corollary 2.3.
Remark 2.
Let $M$ be an $n\times n$ matrix over a field. For any subset $S\subseteq\{1,2,\ldots,n\}$, let $M[S,S]$ denote the submatrix of $M$ with rows and columns indexed by $S$. The following proposition is well known. (For example, see equation (1.2.13) in Horn and Johnson’s book  [Horn2012].)
Proposition 2.
The coefficient $a_{k}$ of $x^{k}$ in $\chi(M)$ satisfies
$a_{n-k}=(-1)^{k}\sum_{\begin{subarray}{c}\#S=k\\ {S\subseteq\{1,2,\ldots,n\}}\end{subarray}}|M[S,S]|.$
But the number of such subsets $S$ is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau,p}})$ somehow force $|J_{n}(\overline{j_{\tau,p}})[S,S]|=(-1)^{k}k!\tau(p_{k})$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^3] See ‘too good.ipynb’in [Bre25]. Instead, there are (we hope) systematic cancellations coming from those symmetries. So far, we do not understand these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau,p}})$.
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau,p}})=\Pi_{n}^{(1)}(x)$ (say.) As we said, in our earlier article [Bre25a] we described regularities in the zero sets of $\Pi_{n}^{(1)}(x)$. Specifically, let $V_{n}^{(1)}$ be the zeros of $\Pi_{n}^{(1)}(x)$ in $\mathbb{C}$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ form a nearly-periodic oscillating curve.
Figure 1. Minimum moduli of the roots of the $\chi_{\Pi^{(1)}_{n}}$
and their logarithms. (Deformed matrices with $c=1$.)
The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the eigenvalues of $\Pi_{n}(x)$ (on which Lehmer’s question hang) to the origin may behave in a regular way, but the details are not yet understood by us because (among other reasons) our data set is very limited and because we do not know the details of the relationship of the eigenvalues of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$. (The corresponding plots for $c=0$ exhibit similar behavior. [^4] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].)
The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 behaves regularly: the abscissas $n$ for a certain subset of the integers in $\{2,3,...,400\}$ of the points on the lower envelope ((corresponding to characteristic polynomial $\Pi_{n}(x)$) for the same subset), after the first three points are dropped, is
5 9 13 17 21 25 29 33 37
41 46 50 54 58 62 66 70 74
78 83 87 91 95 99 103 107 111
115 120 124 128 132 136 140 144 148
152 157 161 165 169 173 177 181 185
189 194 198 202 206 210 214 218 222
226 231 235 239 243 247 251 255 259
263 268 272 276 280 284 288 292 296
300 305 309 313 317 321 325 329 333
337 342 346 350 354 358 362 366 370
374 379 383 387 391 395.
The successive difference of these abscissas are
4 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4.
where the value $5$ appears at the positions congruent to $1$ modulo $9$ (counting from $1$ but omitting $1$.)
Another description of this data is the following. When the abscissas $\{5,9,...,395\}$ are replaced by their residues modulo $4$, the result is
1 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3 3 3 3
3 0 0 0 0 0 0 0 0
0 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3 3 3 3
3 0 0 0 0 0 0 0 0
0 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3.
The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others. [^5] We are using data from the lower envelope of the logarithms of the minimum moduli. (The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25]. These observations suggest the residues of the sequence $\{\tau(p_{n})\}_{n=1,2,...}$ might be interesting to examine. (We have not done that yet.)
For contrast, we reproduce in Figure 2 another plot from our article [Bre25a]: the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$. We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$. The irregularities visible in Figure 2 do not seem promising to us in that regard. This is an obstacle to using these plots to look at Lehmer’s question, obviously, because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the map $n\mapsto\tau(p_{n})$ in a straightforward way. On the other hand, the issue may not be hopeless, in view of the way they were constructed from the values of that map.
Figure 2. Minimum moduli of the roots of undeformed $\chi_{\Pi_{n}}$
We studied the polynomials $\Pi_{n}^{(0)}(x))$. In the range n=1,2,,200n=1,2,...,200, they are monic of degree $n$, their $x^{n-1}$ coefficients are each zero, and they satisfy the following descriptions. [^6] See ‘c=0 deformed vs undeformed tauprime26oct25no3.ipynb’ in [Bre25]. Writing $\Pi_{n}^{(c)}(x)=\sum_{k=0}^{n}a_{c,n,k}x^{k}$, our observations satisfy the following equations:
(4) $a_{0,n,n-2}=-2^{3}*3*\binom{n}{2}$
(5) $a_{0,n,n-3}=2^{4}*3^{2}*\binom{n}{3}$
(6) $a_{0,n,n-4}=2^{2}*3^{3}*1061*\binom{n}{4}$
(7) $a_{0,n,n-5}=-2^{8}*3*16249*\binom{n}{5}$
(8) $a_{0,n,n-6}=2^{6}*3^{2}*5*13*32003*\binom{n}{6}$
(9) $a_{0,n,n-7}=-2^{6}*3^{4}*13*29*53*173*\binom{n}{7}$
(10) $a_{0,n,n-8}=-2^{4}*3^{3}*7*3181*2139307*\binom{n}{8}$
and [^7] See ’c=1&c=1vs0 deformed tauprime29oct25’ and ‘c=0,1 deformed tauprime2nov25.ipynb’ in [Bre25].
(11) $a_{1,n,n-2}=-23*\binom{n}{2}$
(12) $a_{1,n,n-3}=5*43*\binom{n}{3}$
(13) $a_{1,n,n-4}=7*16267*\binom{n}{4}$
(14) $a_{1,n,n-5}=-29*311*1447*\binom{n}{5}$
(15) $a_{1,n,n-6}=7*182111899*\binom{n}{6}$
(16) $a_{1,n,n-7}=-26573015431*\binom{n}{7}$
(17) $a_{1,n,n-8}=-5*53*17827*4318469*\binom{n}{8}$
In spite of appearances, the numerical coefficients of the binomial coefficients on the right side of the equations (11) - (17) are not always square-free. Apparently, these lists might be extended indefinitely.
Of course, the characteristic functions $\Pi_{n}^{(c)}(x)$ may also be written down in complete generality using Theorem 2.3.
The coefficients for $c=0$ and $c=1$ appear to satisfy another, related, description. [^8] See ‘c=0 deformed tauprime28oct25no3.ipynb’ in [Bre25]. For k=2,,200k=2,...,200, at least,
$a_{c,n,n-k}=a_{c,k-1,0}\cdot\binom{n}{k}=\left|J_{k-1}^{(c)}(\overline{j_{\tau,p}})\right|\cdot\binom{n}{k}$.

### v2

@@
 Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] .
-For this reason, as an example in a recent article [Bre25a] about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{1}=2$ and, more generally, $p_{n}$ denotes the $n^{th}$ prime.
-We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit rough regularities: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in an approximately periodic fashion with $n$.
-For obvious reasons, we do not expect that these regularities are sustained without modification throughout the sequence $\{\tau(p_{n})\}$ or any of the other sequences we have examined.
-2. Lemmas and definitions
-To introduce the matrices that we will associate to the members of the sequence $\{\tau(p_{n})\}_{n\geq 1}$, we repeat some definitions and lemmas from our earlier article [Bre25a].
-Although we supplied our own proofs of the lemmas (omitted here), they are well known in the theory of symmetric functions [M].
+For this reason, we made computer calculations about the function on positive integers
+$\tau_{p}:n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime.
+Using matrix machinery reproduced in Lemmas 2.1 and 2.2 below, within the range of our observations, the eigenvalues of the matrices $H_{n}$ and $J_{n}$ attached to this function by the two lemmas behave as follows: the minimum modulus among the eigenvalues for each of the two matrices associated to a particular $\tau_{p}(n)$ appear to oscillate with $n$.
+The connection to Lehmer’s question is that the associated matrix $H_{n}$ has a zero eigenvalue if and only the associated matrix $J_{n}$ does too, and, if they do, $\tau_{p}(n)=0$.
+We show plots of the oscillations in Figure 1 and Figure 3 below.
+We also describe some pleasant patterns among the abscissas of the points on their lower envelopes that seem unlikely to persist unaltered: they seem to suggest that the effect of Ramanujan’s function on the prime sequence is to tune somehow the noise in it, and we do not know how that might work.
+We have looked at similar compositions of tau with other functions and, for them, this “tuning” is not so clear-cut.
+2. Lemmas
+In the context of the theory of symmetric functions, the lemmas in this section are well known.
+They appear for example, in MacDonald’s book [M]), but the proofs are not easy to locate.
+(In fact, we have not located them.) We have deposited home-made proofs of these statements in our depository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”.
 In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@
 $t\frac{d}{dt}H(t)=H(t)J(t).$
-With $f_{0}=1$, let $\overline{f}$ denote the sequence $f_{0},f_{1},f_{2},$.... Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
+With $f_{0}$ possibly defined as equal to one,, let $\overline{f}$ denote the sequence $\{f_{n}\}_{n\in\mathbb{Z}^{+}}$. Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
 $\left(\begin{matrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-1}&j_{n-2}&\cdots&j_{1}&-n+1\\ j_{n}&j_{n-1}&j_{n-2}&\cdots&j_{1}\\ \end{matrix}\right)$
@@
 for $n\geq 1$.
-We define deformations $J^{(c)}_{n}(\overline{j})$ of the $J_{n}(\overline{j})$ as the matrices
-$\left(\begin{matrix}c&-1&0&\cdots&0\\ j_{1}&c&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-2}&j_{n-3}&\cdots&c&-n+1\\ j_{n-1}&j_{n-2}&j_{n-3}&\cdots&c\\ \end{matrix}\right).$
-Remark 1.
-At present we have no helpful insight into the theory of these deformed matrices except to say that (obviously) $J^{(c)}_{n}(\overline{j})=J_{n}\left(\overline{j}^{(c)}\right)$ where $\overline{j}^{(c)}=\{c,j_{1},j_{2},$...\}.
-Now we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999]. Let $A_{n}=$
+Next we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999].
+Proposition 1.
+(Vein and Dale, Theorem 4.23) Let $A_{n}=$
 $\begin{pmatrix}a_{1}&-1&0&\cdots&0\\ a_{2}&a_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n-1}&a_{n-2}&\cdots&a_{1}&-(n{-}1)\\ a_{n}&a_{n-1}&\cdots&a_{2}&a_{1}\end{pmatrix}^{T},$
 where the $T$ operator is matrix transpose. Let $B_{n}(x)=|A_{n}-xI_{n}|$, so that $B_{n}(x)=(-1)^{n}\chi_{A_{n}}(x)$. Then
-Proposition 1.
-(Vein and Dale, Theorem 4.23)
 $B_{n}(x)=\sum_{r=0}^{n}\binom{n}{r}|A_{r}|x^{n-r}.$
 Theorem 2.3.
- [^1] Here we are merely codifying a method shown to us by J. López-Bonilla. (See Corollary 3.3 in our previous paper [Bre25a].)
+ [^2] Here we are merely codifying an argument shown to us by J. López-Bonilla.
 With $h$ and $j$ as in the above lemmas,
@@
 $\begin{pmatrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ j_{n-1}&a_{n-2}&\cdots&j_{1}&-(n{-}1)\\ j_{n}&j_{n-1}&\cdots&j_{2}&j_{1}\end{pmatrix}$.
-Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. Then $C_{n}(x)=$
+Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. By Proposition 1, $C_{n}(x)=$
 $\sum_{r=0}^{n}\binom{n}{r}|(J_{r}(\overline{j})|(-1)^{n-r}x^{n-r}=$
@@
 Simplifying, $\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
-3. Empirical observations about the sequence $\{\tau(p_{n})\}_{n\geq 1}$
-3.1. Characteristic functions of the undeformed matrices representing the function $n\mapsto\tau(p_{n})$
-Let $h_{\tau,p}(0)=1$ and $h_{\tau,p}(n)=\tau(p_{n})$ for positive integers $n$.
-Let $j_{\tau,p}(n)$ and $h_{\tau,p}(n)$ satisfy equation (D) together.
-Clearly, the values of $j_{\tau,p}(n)$ can be computed from equation (D1).
-Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau,p}}))=|xI-J_{n}(\overline{j_{\tau,p}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau,p}})$.
-It is clear that the following statements are equivalent: (a) $\tau(p_{n})=0$, (b) $|J_{n}(\overline{j_{\tau,p}})|=0$, and (c) the polynomial $\chi(J_{n}(\overline{j_{\tau,p}}))$ takes zero to zero.
-In the range of our observations, the degree of $\Pi_{n}(x)$ is $n$. Writing $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k},$ [^2] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25].
-(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau(p_{k})$
-(also in the range of our observations.) Equation (3) is immediate after Corollary 2.3.
+3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
+3.1. Characteristic functions of the undeformed matrices of the form $J_{n}(\overline{j_{\tau_{p}}})$ representing the function $\tau_{p}$.
+Let $h_{\tau_{p}}(0)=1$ and $h_{\tau_{p}}(n)=\tau_{p}(n)$ for positive integers $n$.
+Let $j_{\tau_{p}}(n)$ and $h_{\tau_{p}}(n)$ together satisfy equation (D).
+Given $h_{\tau_{p}}$, the elements of $\overline{j_{\tau_{p}}}=\{j_{\tau_{p}(n)}\}_{n=1,2,...}$ can be computed from equation (D1).
+By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j_{\tau_{p}}})|/n!$.
+Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau_{p}}}))(x)=|xI-J_{n}(\overline{j_{\tau_{p}}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau_{p}}})$.
+The following statements are clearly equivalent: (a) $\tau(p_{n})=0$,
+(b) $|J_{n}(\overline{j_{\tau_{p}}})|=0$, and (c) $\Pi_{n}(0)=0$.
+Writing [^3] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25]. $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k}$, Theorem 2.3 gives
+(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
+In (theoretical) practice, this result would allow us to bypass the construction of the matrices $J_{n}(\overline{j_{\tau_{p}}})$ in our computations, saving time and allowing us to extend our data sets.
+We have not found occasion to exploit this advantage; if a similar result applied to the $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))(x)$, we would have used it; at present, this is not available to us.
 Remark 1.
@@
 $a_{n-k}=(-1)^{k}\sum_{\begin{subarray}{c}\#S=k\\ {S\subseteq\{1,2,\ldots,n\}}\end{subarray}}|M[S,S]|.$
-But the number of such subsets $S$ is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau,p}})$ somehow force $|J_{n}(\overline{j_{\tau,p}})[S,S]|=(-1)^{k}k!\tau(p_{k})$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$.
+The number of terms in the sum is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$.
 But this turns out to be too good to be true.
 [^4] See ‘too good.ipynb’in [Bre25].
-Instead, there are (we hope) systematic cancellations coming from those symmetries.
-So far, we do not understand these cancellations.
-3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau,p}})$.
-Let $\chi(J_{n}^{(1)}(\overline{j_{\tau,p}})=\Pi_{n}^{(1)}(x)$ (say.) As we said, in our earlier article [Bre25a] we described regularities in the zero sets of $\Pi_{n}^{(1)}(x)$.
-Specifically, let $V_{n}^{(1)}$ be the zeros of $\Pi_{n}^{(1)}(x)$ in $\mathbb{C}$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ form a nearly-periodic oscillating curve.
-Figure 1. Minimum moduli of the roots of the $\chi_{\Pi^{(1)}_{n}}$
-and their logarithms. (Deformed matrices with $c=1$.)
-The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the eigenvalues of $\Pi_{n}(x)$ (on which Lehmer’s question hang) to the origin may behave in a regular way, but the details are not yet understood by us because (among other reasons) our data set is very limited and because we do not know the details of the relationship of the eigenvalues of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$.
-(The corresponding plots for $c=0$ exhibit similar behavior.
-[^4] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].)
-The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 behaves regularly: the abscissas $n$ for a certain subset of the integers in $\{2,3,...,400\}$ of the points on the lower envelope ((corresponding to characteristic polynomial $\Pi_{n}(x)$) for the same subset), after the first three points are dropped, is
-5 9 13 17 21 25 29 33 37
-41 46 50 54 58 62 66 70 74
-78 83 87 91 95 99 103 107 111
-115 120 124 128 132 136 140 144 148
-152 157 161 165 169 173 177 181 185
-189 194 198 202 206 210 214 218 222
-226 231 235 239 243 247 251 255 259
-263 268 272 276 280 284 288 292 296
-300 305 309 313 317 321 325 329 333
-337 342 346 350 354 358 362 366 370
-374 379 383 387 391 395.
-The successive difference of these abscissas are
-4 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4 4 4 4 4 4
-5 4 4 4.
-where the value $5$ appears at the positions congruent to $1$ modulo $9$ (counting from $1$ but omitting $1$.)
-Another description of this data is the following. When the abscissas $\{5,9,...,395\}$ are replaced by their residues modulo $4$, the result is
+Instead, there must be systematic cancellations coming from those symmetries.
+So far, we have not understood these cancellations.
+3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau_{p}}})$.
+[…]
+Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here.
+[^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve.
+Figure 1. Minimum moduli of the roots of the $\Pi^{(1)}_{n}(x)$
+and their logarithms.
+Figure 2. Minimum moduli of the roots of undeformed $\Pi_{n}(x)$.
+The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the roots of $\Pi_{n}^{(1)}(x)$ to the origin may behave in a predictable way, but the details are not clear to us because (among other reasons) our data is limited.
+Also, we do not know the details of the relationship of the roots of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$, so the connection, if any, to Lehmer’s question is also unclear.
+Furthermore, we have no statement like equation (3) for the $\Pi_{n}^{(c)}(x)$, because we do not have an explicit description of the sequence $\{h_{n}\}$ represented in Lemma 2.2 by the matrices of which the $\Pi_{n}^{(c)}(x)$ are the characteristic polynomials.
+As we said above, such a description would permit us to sidestep the construction of the matrices we currently use to find these polynomials and the resulting time savings would in practice allow us to collect more data.
+More data would be welcome in view of the oscillations we are trying to check.
+The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 appear to behave regularly: the abscissas of the points on the lower envelope form a set $\{5,9,...,395\}$, which, when they are replaced by their residues modulo $4$, make Table 1.
 1 1 1 1 1 1 1 1 1
@@
 1 2 2 2 2 2 2 2 2
-2 3 3 3 3 3.
+2 3 3 3 3 3
+Table 1. Residues modulo 4 of abscissas from the lower envelope of Figure 1.
 The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others.
@@
 (The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25].
-These observations suggest the residues of the sequence $\{\tau(p_{n})\}_{n=1,2,...}$ might be interesting to examine.
-(We have not done that yet.)
-For contrast, we reproduce in Figure 2 another plot from our article [Bre25a]: the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$.
+For contrast, we reproduce in Figure 2 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$.
 We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$.
 The irregularities visible in Figure 2 do not seem promising to us in that regard.
-This is an obstacle to using these plots to look at Lehmer’s question, obviously, because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the map $n\mapsto\tau(p_{n})$ in a straightforward way.
-On the other hand, the issue may not be hopeless, in view of the way they were constructed from the values of that map.
-Figure 2. Minimum moduli of the roots of undeformed $\chi_{\Pi_{n}}$
-We studied the polynomials $\Pi_{n}^{(0)}(x))$.
-In the range n=1,2,,200n=1,2,...,200, they are monic of degree $n$, their $x^{n-1}$ coefficients are each zero, and they satisfy the following descriptions.
-[^6] See ‘c=0 deformed vs undeformed tauprime26oct25no3.ipynb’ in [Bre25].
-Writing $\Pi_{n}^{(c)}(x)=\sum_{k=0}^{n}a_{c,n,k}x^{k}$, our observations satisfy the following equations:
-(4) $a_{0,n,n-2}=-2^{3}*3*\binom{n}{2}$
-(5) $a_{0,n,n-3}=2^{4}*3^{2}*\binom{n}{3}$
-(6) $a_{0,n,n-4}=2^{2}*3^{3}*1061*\binom{n}{4}$
-(7) $a_{0,n,n-5}=-2^{8}*3*16249*\binom{n}{5}$
-(8) $a_{0,n,n-6}=2^{6}*3^{2}*5*13*32003*\binom{n}{6}$
-(9) $a_{0,n,n-7}=-2^{6}*3^{4}*13*29*53*173*\binom{n}{7}$
-(10) $a_{0,n,n-8}=-2^{4}*3^{3}*7*3181*2139307*\binom{n}{8}$
-and [^7] See ’c=1&c=1vs0 deformed tauprime29oct25’ and ‘c=0,1 deformed tauprime2nov25.ipynb’ in [Bre25].
-(11) $a_{1,n,n-2}=-23*\binom{n}{2}$
-(12) $a_{1,n,n-3}=5*43*\binom{n}{3}$
-(13) $a_{1,n,n-4}=7*16267*\binom{n}{4}$
-(14) $a_{1,n,n-5}=-29*311*1447*\binom{n}{5}$
-(15) $a_{1,n,n-6}=7*182111899*\binom{n}{6}$
-(16) $a_{1,n,n-7}=-26573015431*\binom{n}{7}$
-(17) $a_{1,n,n-8}=-5*53*17827*4318469*\binom{n}{8}$
-In spite of appearances, the numerical coefficients of the binomial coefficients on the right side of the equations (11) - (17) are not always square-free.
-Apparently, these lists might be extended indefinitely.
-Of course, the characteristic functions $\Pi_{n}^{(c)}(x)$ may also be written down in complete generality using Theorem 2.3.
-The coefficients for $c=0$ and $c=1$ appear to satisfy another, related, description. [^8] See ‘c=0 deformed tauprime28oct25no3.ipynb’ in [Bre25]. For k=2,,200k=2,...,200, at least,
-$a_{c,n,n-k}=a_{c,k-1,0}\cdot\binom{n}{k}=\left|J_{k-1}^{(c)}(\overline{j_{\tau,p}})\right|\cdot\binom{n}{k}$.
+This is an obstacle to using these plots to look at Lehmer’s question because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the function $\tau_{p}$ in a straightforward way.
+3.3. Characteristic functions of deformed matrices $H_{n}^{(1)}(\overline{h_{\tau_{p}}})$.
+ [^7] See our notebook “tauprime=j H matrix4dec25no2.ipynb” in [Bre25].
+Given the sequence $\{\tau_{p}(n)\}_{n=1,2,...}=$ (say) $\{j_{\tau_{p}}(n)\}_{n=1,2,...}=\overline{j_{\tau_{p}}}$, we
+construct a sequence $\overline{h_{\tau_{p}}}=\{h_{\tau_{p}}(n)\}_{n=1,2,...}$ by solving the convolution (D1). By Lemma 2.2, $\tau_{p}(n)=(-1)^{n+1}|H_{n}(\overline{h})|$. Then we construct matrices
+[…]
+Figure 3. Minimum moduli and their logarithms for zeros of the $\Psi^{(1)}_{n}(x)$.
+The table of modulo $4$ residues for the abscissas of the lower envelopes of the plots in Figure 3 is not as regular as the corresponding table for Figure 1; the lengths of the strings of constant residue vary:
+0 0 2 2 3 3 3 3 3
+0 0 0 0 0 1 1 1 1
+1 1 2 2 2 2 2 3 3
+3 3 3 3 0 0 0 0 0
+1 1 1 1 1 1 2 2 2
+2 2 2 2 3 3 3
+Table 2. Residues modulo 4 of abscissas from thelower envelope of Figure 3.

### v3

@@
-Ramanujan’s function on small primes
+Ramanujan functions on small primes
 Abstract.
-Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number.
-For this reason, as an example in a recent article about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime.
-We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit regularities of the following kind: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in a nearly periodic fashion with $n$.
+Lehmer asked whether Ramanujan’s tau function has any zeros and showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number.
+Certain matrix identities originating in the theory of symmetric functions express members of arithmetic sequences as determinants.
+We study empirically the eigenvalues of the matrices underlying them with the goal of estimating how close the eigenvalues of the matrix associated to the number $\tau(p)$ for $p$ prime get to zero, because (with our set-up) a zero eigenvalue would entail that $\tau(p)$ vanishes.
+We also studied corresponding eigenvalues of several cusp forms coming from elliptic curves.
 Key words and phrases:
@@
 05A17, 05C62, 11F11, 11P83
-11 December 2025
+10 March 2026 no4
 1. Introduction
 Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47].
-For this reason, we made computer calculations about the function on positive integers
-$\tau_{p}:n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime.
-Using matrix machinery reproduced in Lemmas 2.1 and 2.2 below, within the range of our observations, the eigenvalues of the matrices $H_{n}$ and $J_{n}$ attached to this function by the two lemmas behave as follows: the minimum modulus among the eigenvalues for each of the two matrices associated to a particular $\tau_{p}(n)$ appear to oscillate with $n$.
-The connection to Lehmer’s question is that the associated matrix $H_{n}$ has a zero eigenvalue if and only the associated matrix $J_{n}$ does too, and, if they do, $\tau_{p}(n)=0$.
-We show plots of the oscillations in Figure 1 and Figure 3 below.
-We also describe some pleasant patterns among the abscissas of the points on their lower envelopes that seem unlikely to persist unaltered: they seem to suggest that the effect of Ramanujan’s function on the prime sequence is to tune somehow the noise in it, and we do not know how that might work.
-We have looked at similar compositions of tau with other functions and, for them, this “tuning” is not so clear-cut.
+Matrix identities ([M], Lemmas 2.1 and 2.2 below) govern certain additive convolution identities.
+We study eigenvalues of these matrices empirically, especially eigenvalues associated to the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime, because (with our set-up) a zero eigenvalue would entail that $\tau(p_{n})$ vanishes.
+After a procedure we will call “deformation”, the minimum modulus among the eigenvalues for a given $n$ appears to oscillate in a nearly periodic fashion with $n$.
+(We also saw this phenomenon in the sequence $\{\tau(p_{n+1}-p_{n})\}_{n=1,2,...}$; we omit it here.) We have also observed oscillatory behavior in several newforms coming from elliptic curves in virtue of the Modularity Theorem.
+To encompass those observations, given the appropriate uniformizing function $q_{*}(z)=q_{*}$ (say) from the upper half plane $\mathbb{H}$ to the unit disk for a cusp form $f(z)=\sum_{n}a(n)q_{*}^{n}$, we have referred in our title to the function $n\mapsto a(n)$ as a Ramanujan function.
 2. Lemmas
-In the context of the theory of symmetric functions, the lemmas in this section are well known.
-They appear for example, in MacDonald’s book [M]), but the proofs are not easy to locate.
-(In fact, we have not located them.) We have deposited home-made proofs of these statements in our depository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”.
+The lemmas in this section are well known.
+They appear, for example, in MacDonald’s book [M]), but the proofs are not easy to locate.
+(In fact, we have not located them.) We have placed home-made proofs of these statements in our repository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”.
 In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@
 $B_{n}(x)=\sum_{r=0}^{n}\binom{n}{r}|A_{r}|x^{n-r}.$
+The corollary is the following
 Theorem 2.3.
@@
 Simplifying, $\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
-3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
-3.1. Characteristic functions of the undeformed matrices of the form $J_{n}(\overline{j_{\tau_{p}}})$ representing the function $\tau_{p}$.
-Let $h_{\tau_{p}}(0)=1$ and $h_{\tau_{p}}(n)=\tau_{p}(n)$ for positive integers $n$.
-Let $j_{\tau_{p}}(n)$ and $h_{\tau_{p}}(n)$ together satisfy equation (D).
-Given $h_{\tau_{p}}$, the elements of $\overline{j_{\tau_{p}}}=\{j_{\tau_{p}(n)}\}_{n=1,2,...}$ can be computed from equation (D1).
-By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j_{\tau_{p}}})|/n!$.
-Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau_{p}}}))(x)=|xI-J_{n}(\overline{j_{\tau_{p}}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau_{p}}})$.
-The following statements are clearly equivalent: (a) $\tau(p_{n})=0$,
-(b) $|J_{n}(\overline{j_{\tau_{p}}})|=0$, and (c) $\Pi_{n}(0)=0$.
-Writing [^3] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25]. $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k}$, Theorem 2.3 gives
-(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
-In (theoretical) practice, this result would allow us to bypass the construction of the matrices $J_{n}(\overline{j_{\tau_{p}}})$ in our computations, saving time and allowing us to extend our data sets.
-We have not found occasion to exploit this advantage; if a similar result applied to the $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))(x)$, we would have used it; at present, this is not available to us.
-Remark 1.
-Let $M$ be an $n\times n$ matrix over a field.
+Now let $M$ be an $n\times n$ matrix over a field.
 For any subset $S\subseteq\{1,2,\ldots,n\}$, let $M[S,S]$ denote the submatrix of $M$ with rows and columns indexed by $S$.
@@
 $a_{n-k}=(-1)^{k}\sum_{\begin{subarray}{c}\#S=k\\ {S\subseteq\{1,2,\ldots,n\}}\end{subarray}}|M[S,S]|.$
-The number of terms in the sum is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$.
+3. The sequences $\{p_{n}\}_{n\geq 1}$ and $\{\tau(p_{n})\}_{n\geq 1}$
+[…]
+We will freely abuse the notation by employing the $H,h,J$, and $j$ symbols to denote objects in a way that depends on the context without, we hope, confusing anyone.
+3.1. The sequence $\{\tau(p)\}_{p\text{ prime}}$ and its developments
+Let $p_{n}$ denote the $n^{th}$ prime, $h(0)=1$ and $h(n)=\tau(p_{n})=$ (say) $\tau_{p}(n)$ for positive integers $n$.
+Let $j(n)$ and $h(n)$ together satisfy equation (D).
+Given $h$, the elements of $\overline{j}=\{j\}_{n=1,2,...}$ can be computed from equation (D1).
+By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j})|/n!$.
+Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j}))(x)=|xI-J_{n}(\overline{j})|$ be the characteristic polynomial of $J_{n}(\overline{j})$.
+Plots 3 and 4 show the behavior of the minimum modulus among the eigenvalues of $\Pi^{(1)}_{n}$ and $\Pi_{n}$ respectively.
+They are relevant to Lehmer’s question because the following statements are equivalent:
+(a) $\tau(p_{n})=0$,
+(b) $|J_{n}(\overline{j})|=0$, and
+(c) $\Pi_{n}(0)=0$.
+Remark 1.
+Writing [^3] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25]. $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k}$, Theorem 2.3 gives
+(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
+This result allows bypass of the construction of the matrices $J_{n}(\overline{j})$, in computations, saving time and allowing the extension our data sets.
+We have not exploited this advantage; if a similar result applied to the deformations $\chi(J_{n}^{(c)}(\overline{j}))(x)$ (see section 3.2) we would have used it but, at present, this is not available to us.
+Remark 2.
+The number of terms in the sum in Proposition 2 is $\binom{n}{k}$, so the simplest way to reconcile Proposition 2 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$.
 But this turns out to be too good to be true.
@@
 So far, we have not understood these cancellations.
-3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau_{p}}})$.
-[…]
+3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j})$ representing the sequence $\{\tau(p_{n})\}_{n\geq 1}$
 Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here.
 [^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve.
-Figure 1. Minimum moduli of the roots of the $\Pi^{(1)}_{n}(x)$
-and their logarithms.
-Figure 2. Minimum moduli of the roots of undeformed $\Pi_{n}(x)$.
+As a control, we repeat the calculation for $h(n)=\tau(p_{n}+1)$.
+Figure 2 shows no oscillatory behavior for this choice of $h$, indicating that the choice of prime indices per se was decisive.
 The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the roots of $\Pi_{n}^{(1)}(x)$ to the origin may behave in a predictable way, but the details are not clear to us because (among other reasons) our data is limited.
@@
 The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 3 appear to behave regularly: the abscissas of the points on the lower envelope form a set $\{5,9,...,395\}$, which, when they are replaced by their residues modulo $4$, make Table 1.
+The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others.
+[^6] We are using data from the lower envelope of the logarithms of the minimum moduli.
+(The lower plot in Figure 3.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25].
+(For contrast, we reproduce in Figure 4 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$.
+We have not attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$.)
+Fourier analyses were done by the Anthropic AI “Claude” [Anthropic2026]. (We will supply more details and also put Claude’s analyses verbatim in our depository in later drafts.)
+Next, with $h_{n}=\tau_{p}(n)$, we used equation (D1) to compute the sequence $\overline{j_{\tau_{p}}}$.
+We then “deformed” it by prepending $c$ to $\overline{j_{\tau_{p}}}$.
+We denote this deformation as $\overline{j^{(c)}_{\tau_{p}}}$.
+Then, once more applying equation (D1), we compute a sequence $\overline{h^{(c)}_{\tau_{p}}}$ from $\overline{j^{(c)}_{\tau_{p}}}$.
+We formed the list of pairs $(c,n!h^{(c)}(n))$.
+(These are just the values of the determinants from Lemma 2.2 in this situation; but it is a situation in which we can bypass calculating the determinants and, by that expedient, achieve significant time savings.)
+It seemed likely that these pairs would describe a polynomial in $c$, $D_{n}(c)$ (say.) We found by Lagrange interpolation this is so within the range of our observations, and that the degree of $D_{n}(c)$ appears to be $n$.
+We found the minimum modulus among the roots of $D_{n}(c)$ and plotted these minima against $n$.
+In the case $c\in\mathbb{Z}^{\geq 0}$, the result was the same kind of oscillating wave form, as far as we could see (literally.) [^7] Figure 5, Table 7, and notebook ‘tauprime many c’s 18feb26’ in [Bre25].
+The reader will notice that we ran the routines twice, at 64-bit and 100-bit precision; the results were indistinguishable.
+The Fourier analysis of this waveform is described in Table 2.
+4. Oscillatory behavior in newforms from elliptic curves
+Because the cusp form $\Delta$ is the generating function of the multiplicative Ramanujan function, which displays oscillatory behavior in the range of our observations, we searched a source of cusp forms that are generating functions of multiplicative functions, and found similar oscillatory behavior: Cremona’s database [LMFDB] of cusp forms $\sum_{n}a(n)q_{*}^{n}$ for an appropriate uniformizing parameter $q_{*}(z)$ associated to elliptic curves in virtue of the Modularity Theorem.
+These cusp forms have weight two at various levels.
+[^8] SageMath notebooks corresponding to Table 4 are in [Bre25].
+In the limited range of our observations, some deformed curves (in the sense we have been using that term) show oscillatory waveforms, and some that do not, do show oscillatory behavior when the sequences $\{h\}$ satisfy $h(n)=a(1+p_{n})$.
+[^9] We had thought, mistakenly, that our code was checking the coefficients $a(p_{n})$.
+To assure ourselves that our present interpretation of the code is the correct one, we recomputed some of the plots with code that checks it explicitly.
+For example, see the notebooks “curve 11a1 $h(n)=a(p_{n}+1)$ 8mar26” and “curve 14a1 $h(n)=a(p_{n}+1)$ 9mar26” in [Bre25].
+Table 3 summarizes Claude’s findings.
+[^10] The analysis of the curve $\tau(p_{n}),c=0$ was carried out in the notebook ‘deformed c=0 primetau3feb26’ and that of the curve $\tau(p_{n}),c=1$, was done in the notebook ‘tauprime=h logarithmic envelopes prec100 n400 1dec25’, both stored in [Bre25].
+The curve labels indicate characteristics of the associated elliptic curves.
+(See the explanation in section 6.1 below of Table 8, column A.) The column labeled “cc” indicates the deformation parameter in the sense described above.
+The column labeled “Indices” distinguishes between matrices for $h(n)=a_{n}$ and $h(n)=a_{p_{n}}$.
+More detail will follow in subsequent drafts of this article.
+5. Questions
+We have a dilemma.
+The plots for the undeformed matrices are a mess, therefore, unpromising as sources of insight about the distances of their eigenvalues from the origin, but their determinants represent the underlying $h$ sequences in a straightforward way.
+The plots for the deformed matrices seem more promising.
+It is possible that eventually we will be able to use them to formulate some guesses bounding the plotted minimum modulus eigenvalues away from zero.
+But, so far, we cannot use such conjectures to bound their plotted minimal modulus eigenvalues in the undeformed setting away from zero, and that is what we need.
+Beyond the matter of making guesses, we do not know why the oscillations we see in the deformed setting occur, why they are absent in the undeformed setting, or very much about how the two settings are related.
+Other questions are: Are the members of $\overline{h^{(c)}_{\tau_{p}}}$ Fourier coefficients of a modular form?
+What conditions on $h(n)$ produce oscillatory behavior among the $h^{(c)}$?
+To help with comparisons between the presence or absence of oscillatory behavior and characteristics of the objects that do or do not manifest it, we list cusp forms and associated invariants computed using SageMath native functions [sagemath] (Tables 8 - 11.) The tables are explained in the next section.
+[^11] See ‘cusp form gp invariants 24feb26’ in [Bre25].
+Tables 8 - 11 and the key to them in the next section were created to our specifications by the AI “Claude”[Anthropic2026], and edited by the author.
+6. Key to Tables 8 - 11
+The tables enumerate several invariants for Ramanujan’s $\Delta$ function and elliptic curves over $\mathbb{Q}$ with conductor $N<54$ and unique Fourier expansions in the Cremona database [6].
+(Thus there is only one curve labeled NXm, as explained below, for each NX.) The key omits discussion of the two rightmost columns in each table.
+These are yes or no responses to the questions stated in the table captions.
+At the time of the present draft, they are entirely subjective.
+6.1. Table 8: Basic Arithmetic Invariants
+Column A: Curve Label
+The Cremona label uniquely identifies an elliptic curve over $\mathbb{Q}$ in the form $NXm$, where $N$ is the conductor, $X$ is a letter indicating the isogeny class (a, b, c, …), and $m$ is a number ordering curves within the isogeny class by discriminant.
+Isogenous curves have the same associated cusp form (“newform”).
+For Ramanujan’s $\Delta$ function, we use the label “Delta”.
+Code:
+label = E_cremona.label() # Line 267
+# For Delta: label = "Delta" # Line 176
+References: The Cremona labeling scheme is documented in [5]. Isogeny class structure is discussed in [13, Chapter VII].
+Column B: Minimal Weierstrass Form
+The minimal (global) Weierstrass model of the curve, reported as the coefficient vector $a_{1},a_{2},a_{3},a_{4},a_{6}]$ in the equation
+$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}.$
+Code:
+E = EllipticCurve(label) # e.g. EllipticCurve(’37a1’)
+ainvs = list(E.ainvs()) # returns [a1, a2, a3, a4, a6]
+References: Minimal Weierstrass models: [13, Section VII.1]. Uniqueness: [13, Proposition VIII.8.2].
+Column C: Weight
+The weight of the associated modular form. Code:
+weight = 2 # All elliptic curves, Line 286
+# For Delta: weight = 12 # Line 181
+References: Modularity Theorem: [20, 17, 1]. Weight of modular forms: [diamond-shurman, Chapter 5].
+Column D: Torsion Order
+The order $|E(\mathbb{Q})_{\text{tors}}|$ of the finite torsion subgroup of the Mordell-Weil group. Code:
+torsion_order = int(E.torsion_order()) # Line 289
+# For Delta: torsion_order = "N.A." # Line 178
+References: Mazur’s theorem [9]. Computational methods: [3, Algorithm 7.4.12].
+Column E: Torsion Structure
+The group structure of $E(\mathbb{Q})_{\text{tors}}$ as an abelian group, expressed in the form $\mathbb{Z}/n_{1}\mathbb{Z}\times\cdots\times\mathbb{Z}/n_{k}\mathbb{Z}$ where $n_{i}|n_{i+1}$ (invariant factors).
+Code:
+def torsion_structure_string(E):
+T = E.torsion_subgroup()
+inv = T.invariants() # Returns tuple of invariant factors
+if len(inv) == 0:
+return "trivial"
+else:
+return "x".join([f"Z/{n}Z" for n in inv])
+# Line 101-109
+References: The SageMath method torsion_subgroup() implements the algorithm from [18, Section 8.7]. Structure theorem for finitely generated abelian groups: [7, Theorem 5.2]. Computation: [18].
+Column F: Rank
+The rank $r$ of the Mordell-Weil group $E(\mathbb{Q})$. Code:
+try:
+rank = int(E.rank())
+rank_certain = True
+except Exception:
+rank = -1
+rank_certain = False
+# Lines 293-298
+References: SageMath’s E.rank() implements the algorithm of [5], which combines 2-descent with verification using the BSD conjecture for rank 0 and 1 cases.
+Mordell-Weil theorem: [10, 19, 13].
+BSD conjecture: [2].
+Computational methods: [5].
+6.2. Table 9: Mordell-Weil Generators and Galois Representations
+Column A: Curve Label
+Same as column A of Table 8.
+Column B: $j$-Invariant
+The $j$-invariant $j(E)\in\mathbb{Q}$.
+Code:
+E = EllipticCurve(label)
+j = E.j_invariant() # returns an element of $Q$
+References: $j$-invariant and the $b$-, $c$-invariants: [13, Section III.1]. Relation to CM: [14, Appendix A].
+Column C: Generator Canonical Heights
+Néron-Tate height $\hat{h}(P)$.
+Code:
+def generator_heights_string(E):
+try:
+gens = E.gens()
+if len(gens) == 0:
+return "none"
+heights = [float(g.height()) for g in gens] # CORRECTED
+return "|".join([f"{h:.4f}" for h in heights])
+except Exception as ex:
+return f"error:{ex}"
+# Lines 111-119
+References: Canonical height theory: [13, Chapter VIII]. Computational methods: [5, 15].
+Column D: Regulator
+The determinant of the height pairing matrix.
+Code:
+try:
+reg = float(E.regulator())
+except Exception:
+reg = -1.0
+# Lines 306-309
+References: Regulator definition: [13, Section X.6]. BSD conjecture: [2, 8].
+Column E: Galois Representation Surjective for All Primes?
+For a prime $\ell$, this is the Galois representation
+$\rho_{E,\ell}:\text{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\text{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
+arising from the action of the absolute Galois group on the $\ell$-torsion $E[\ell]$.
+Code:
+try:
+rho = E.galois_representation()
+non_surj = list(rho.non_surjective())
+gal_surjective = (len(non_surj) == 0)
+except Exception as ex:
+gal_surjective = "error"
+# Lines 313-317
+References: Serre’s open image theorem: [11, 12]. Computational aspects: [16].
+6.3. Table 10: Galois Exceptional Primes and Modular Data
+Column B: Galois Image Type
+A classification of the image of the mod-$\ell$ Galois representation.
+Possible types: “surjective” ($\text{im}(\rho_{E,\ell})=\text{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ for all $\ell$); “non-surjective at $\ell_{1},\ldots,\ell_{k}]$” (list of exceptional primes); or “CM” (curve has complex multiplication, image contained in normalizer of split/non-split Cartan subgroup).
+The numbers in brackets are the lists of the finite set of primes $\ell$ for which $\rho_{E,\ell}$ is not surjective.
+[11, 12, 16] For non-CM curves over $\mathbb{Q}$, this set is always finite (Serre).
+Code:
+non_surj = list(rho.non_surjective())
+gal_exceptional_primes = "|".join([str(p) for p in non_surj]) \
+if non_surj else "none"
+# Lines 318-319
+def galois_image_type(E):
+try:
+rho = E.galois_representation()
+if E.has_cm():
+return "CM"
+non_surj = rho.non_surjective()
+if len(non_surj) == 0:
+return "surjective"
+else:
+return f"non-surjectiveat{list(non_surj)}"
+except Exception as ex:
+return f"error:{ex}"
+# Lines 121-131
+References: Image classifications: [11, 12, 16] CM theory: [14, Chapter II].
+Column C: Index of $\Gamma_{0}(N)$ in $\text{SL}_{2}(\mathbb{Z})$
+(NN is the conductor of the curve.) Code:
+def X0N_data(N):
+try:
+G = Gamma0(N)
+idx = G.index()
+genus = G.genus()
+cusps = G.ncusps()
+return idx, genus, cusps
+except Exception as ex:
+return "error", "error", "error"
+# Lines 133-141
+References: Congruence subgroups: [diamond-shurman, Chapter 1]. Index formula: [diamond-shurman, Proposition 1.4.4].
+6.4. Table 11: Modular Curves and L-functions
+Column B: Genus of $X_{0}(N)$
+NN is the conductor. $X_{0}(N)$ is the compactification of the quotient $\Gamma_{0}(N)\backslash\mathbb{H}$ where $\mathbb{H}$ is the upper half-plane.
+Code:
+G = Gamma0(N)
+genus = G.genus()
+# Line 136
+References: Modular curves: [diamond-shurman, Chapter 3]. Genus formula: [diamond-shurman, Proposition 3.8.2].
+Column C: Number of Cusps of $X_{0}(N)$
+Code:
+cusps = G.ncusps()
+# Line 137
+References: Cusp counting: [diamond-shurman, Proposition 3.8.3].
+Column D: Has Complex Multiplication?
+Code:
+try:
+is_cm = E.has_cm()
+cm_disc = int(E.cm_discriminant()) if is_cm else 0
+except Exception:
+is_cm = "error"
+cm_disc = "error"
+# Lines 326-330
+References: CM theory: [14, Chapter II], [4]. The 13 CM $j$-invariants: [14, Appendix A].
+Column E: CM Discriminant
+If $E$ does not have CM, this field is 0.
+Code:
+cm_disc = int(E.cm_discriminant()) if is_cm else 0
+# Line 328
+Column F: Root Number
+The root number $w\in\{+1,-1\}$. Code:
+try:
+root_number = int(E.root_number())
+except Exception:
+root_number = "error"
+# Lines 333-336
+References: Functional equation: [18, Chapter 16]. Root number and BSD: [2, 8]. Computational methods: [5, Section 2.14].
+Potential issues
+  1. (1)
+    Rank computation: For curves of large rank ($\geq 2$), E.rank() may be slow or fail.
+The BSD conjecture is used heuristically for rank 0 and 1.
+All curves in this dataset ($N\leq 54$) have rank $\leq 1$, so all rank computations are provably correct.
+  2. (2)
+    Generator computation: For rank 1 curves, E.gens() returns provably correct generators using Cremona’s algorithm. Heights are computed to machine precision ($\approx 10^{-15}$ relative error).
+  3. (3)
+    Galois representations: The method non_surjective() is provably correct for the primes it returns, but could theoretically miss exceptional primes beyond SageMath’s computational limits.
+For $N\leq 54$, all computations are complete.
+7. Tables
 1 1 1 1 1 1 1 1 1
@@
 Table 1. Residues modulo 4 of abscissas from the lower envelope of Figure 2.
-The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others.
-[^6] We are using data from the lower envelope of the logarithms of the minimum moduli.
-(The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25].
-For contrast, we reproduce in Figure 2 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$.
-We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$.
-The irregularities visible in Figure 2 do not seem promising to us in that regard.
-This is an obstacle to using these plots to look at Lehmer’s question because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the function $\tau_{p}$ in a straightforward way.
-3.3. Characteristic functions of deformed matrices $H_{n}^{(1)}(\overline{h_{\tau_{p}}})$.
- [^7] See our notebook “tauprime=j H matrix4dec25no2.ipynb” in [Bre25].
-Given the sequence $\{\tau_{p}(n)\}_{n=1,2,...}=$ (say) $\{j_{\tau_{p}}(n)\}_{n=1,2,...}=\overline{j_{\tau_{p}}}$, we
-construct a sequence $\overline{h_{\tau_{p}}}=\{h_{\tau_{p}}(n)\}_{n=1,2,...}$ by solving the convolution (D1). By Lemma 2.2, $\tau_{p}(n)=(-1)^{n+1}|H_{n}(\overline{h})|$. Then we construct matrices
-[…]
-Figure 3. Minimum moduli and their logarithms for zeros of the $\Psi^{(1)}_{n}(x)$.
-The table of modulo $4$ residues for the abscissas of the lower envelopes of the plots in Figure 3 is not as regular as the corresponding table for Figure 1; the lengths of the strings of constant residue vary:
-0 0 2 2 3 3 3 3 3
-0 0 0 0 0 1 1 1 1
-1 1 2 2 2 2 2 3 3
-3 3 3 3 0 0 0 0 0
-1 1 1 1 1 1 2 2 2
-2 2 2 2 3 3 3
-Table 2. Residues modulo 4 of abscissas from thelower envelope of Figure 3.
+Curve $c$ Period Ratio Power
+$\tau(p_{n})$ 0 4.14 1.000 $1.27\times 10^{4}$
+2.06 2.010 $3.98\times 10^{3}$
+3.68 1.125 $1.34\times 10^{3}$
+37.25 0.111 $1.23\times 10^{3}$
+$\tau(p_{n})$ 1 4.11 1.000 $5.65\times 10^{4}$
+2.06 1.995 $8.60\times 10^{3}$
+3.69 1.114 $5.33\times 10^{3}$
+$\tau(p_{n})$ 2 4.11 1.000 $5.40\times 10^{4}$
+2.06 1.995 $7.63\times 10^{3}$
+3.69 1.114 $4.62\times 10^{3}$
+$\tau(p_{n})$ 3 4.11 1.000 $5.61\times 10^{4}$
+2.06 1.995 $8.37\times 10^{3}$
+3.69 1.114 $5.19\times 10^{3}$
+Table 2.
+Fourier analysis of log-transformed, detrended minimum moduli of eigenvalues of $\chi(|J^{(c)}|)$ for $h(n)=\tau(p_{n})$.
+Data spans $n=2$ to 400 (399 primes) for fixed-$c$ analyses ($c=0,1,2,3$).
+For $c=0$, only 298 primes were analyzed.
+The polynomial secular trend (Table 6) was removed before computing the Fourier transform.
+The Ratio column shows the ratio of the fundamental period to the period in question.
+Curve Period Ratio Power
+11a1 11.76 1.000 $3.83\times 10^{4}$
+5.97 1.970 $4.15\times 10^{3}$
+3.96 2.970 $4.10\times 10^{3}$
+2.96 3.973 $2.40\times 10^{3}$
+14a1 5.63 1.000 $3.41\times 10^{4}$
+2.80 2.011 $4.77\times 10^{3}$
+2.15 2.619 $3.52\times 10^{3}$
+15a1 30.77 1.000 $3.92\times 10^{4}$
+15.38 2.000 $9.26\times 10^{3}$
+10.26 2.999 $5.06\times 10^{3}$
+7.69 4.001 $2.78\times 10^{3}$
+19a 2.84 1.000 $4.09\times 10^{4}$
+3.39 0.838 $4.04\times 10^{3}$
+17.39 0.163 $2.12\times 10^{3}$
+24a1 5.30 1.000 $9.83\times 10^{4}$
+37a1 6.56 1.000 $3.58\times 10^{4}$
+3.31 1.982 $4.83\times 10^{3}$
+2.43 2.700 $4.31\times 10^{3}$
+43a1 7.30 1.000 $4.49\times 10^{4}$
+3.65 2.000 $5.34\times 10^{3}$
+2.43 3.004 $2.61\times 10^{3}$
+53a1 6.35 1.000 $4.00\times 10^{4}$
+2.13 2.981 $5.01\times 10^{3}$
+3.20 1.984 $4.34\times 10^{3}$
+2.68 2.370 $2.47\times 10^{3}$
+Table 3.
+Fourier analysis of minimum moduli for elliptic curves using deformed matrices, with $c=1$, for all indices.
+The Ratio column shows the ratio of the fundamental period to the period in question.
+Curve 24a1 shows a single dominant peak, or a tight cluster at the edge of FFT resolution with high power.
+(We show the doubtful periods in the next table.) The reader will notice that the ratios in rows 15a1 and 43a1 are nearly integers.
+Period Power
+5.2980 $9.83\times 10^{4}$
+5.2805 $8.31\times 10^{4}$
+5.3156 $7.93\times 10^{4}$
+5.2632 $4.69\times 10^{4}$
+5.3333 $4.05\times 10^{4}$
+5.2459 $1.55\times 10^{4}$
+5.3512 $9.91\times 10^{3}$
+5.4054 $7.60\times 10^{3}$
+5.3872 $5.62\times 10^{3}$
+5.4237 $3.78\times 10^{3}$
+5.1948 $3.04\times 10^{3}$
+5.4795 $2.74\times 10^{3}$
+5.2288 $2.49\times 10^{3}$
+5.2117 $2.03\times 10^{3}$
+5.4983 $1.97\times 10^{3}$
+5.1780 $1.66\times 10^{3}$
+5.3691 $1.35\times 10^{3}$
+5.5556 $1.25\times 10^{3}$
+5.4608 $1.21\times 10^{3}$
+5.5749 $1.18\times 10^{3}$
+Table 4. The uncertain cluster of periods for curve 24a1 at all indices. See ‘curve 24a1 all indices 21feb26’ in [Bre25].
+Curve Range Period Ratio Power
+11a1 $n<100$ 2.62 1.000 $1.27\times 10^{3}$
+$n\geq 100$ 2.69 1.000 $3.39\times 10^{4}$
+14a1 All 3.81 1.000 $4.81\times 10^{4}$
+2.11 1.806 $5.92\times 10^{3}$
+4.71 0.809 $3.36\times 10^{3}$
+15a1 All 5.70 1.000 $4.75\times 10^{4}$
+2.85 2.000 $5.13\times 10^{3}$
+5.78 0.986 $3.30\times 10^{3}$
+2.105 2.708 $3.61\times 10^{3}$
+17a1 All 3.81 1.000 $4.52\times 10^{4}$
+2.11 1.806 $4.63\times 10^{3}$
+Table 5.
+Fourier analysis of minimum moduli for elliptic curves using deformed matrices, with $c=1$, for prime indices only.
+The Ratio column shows the ratio of the fundamental period to the period in question.
+For curves 11a1 and 15a1, spurious zeros in the first two data points were removed before analysis.
+Curve 11a1 uniquely exhibits non-stationary behavior with period evolution from 2.62 to 2.69 and a 27-fold power increase at the transition ($n\approx 100$), while other curves show stationary harmonic structure throughout.
+Curves 14a1 and 17a1 share matching top two periods.
+Function $c$ Quadratic Linear Constant
+$\tau(p_{n})$ 0 $5.60\times 10^{-6}$ $2.02\times 10^{-3}$ 2.806
+$\tau(p_{n})$ 1 $1.06\times 10^{-6}$ $5.69\times 10^{-4}$ 2.871
+$\tau(p_{n})$ 2 $2.41\times 10^{-6}$ $1.25\times 10^{-3}$ 2.820
+$\tau(p_{n})$ 3 $4.85\times 10^{-6}$ $1.84\times 10^{-3}$ 2.822
+Table 6.
+Secular trends in log-transformed minimum moduli for data from Table 2.
+The fitted polynomial is $\log(\text{min modulus})\approx a_{2}n^{2}+a_{1}n+a_{0}$, where $n$ is the prime index.
+All cases show small positive linear coefficients, indicating minimum moduli increase slightly with $n$ (moving away from zero).
+The quadratic terms are negligible.
+Function Indices Period Power
+$\tau(p_{n})$ Primes 4.10 $2.29\times 10^{4}$
+Table 7. Global (over $c$) Fourier analysis for $h(n)=\tau_{p}(n)$ and $n=2$ to $300,1\leq c\leq 2n$, degree of $D_{n}(c)=n$.)
+Table 8.
+Basic arithmetic invariants for $\Delta$ and elliptic curves with conductor $\leq 53$.
+Column labels: A = Curve label, B = Minimal Weierstrass form $a_{1},a_{2},a_{3},a_{4},a_{6}]$, C = Weight, D = Torsion order, E = Torsion structure, F = Rank.
+G = Deformed plots with $h(n)=a(n)$ oscillate?
+H = Deformed plots with $h(n)=a(1+p_{n})$ oscillate?
+K = Deformed plots with $h(n)=a(p_{n})$ oscillate?
+The conductor of $\Delta$ is 11 and that of each other cusp form is given by in its label.
+Thus the conductor of 11a1 is $11$.
+A B C D E F G H K
+$\Delta$ 12 N N Y
+11a1 [0,1,1,10,20][0,-1,1,-10,-20] 2 5 $\mathbb{Z}/5\mathbb{Z}$ 0 Y Y N
+14a1 [1,0,1,4,6][1,0,1,4,-6] 2 6 $\mathbb{Z}/6\mathbb{Z}$ 0 Y Y N
+15a1 [1,1,1,10,10][1,1,1,-10,-10] 2 8 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/4\mathbb{Z}$ 0 Y Y
+17a1 [1,1,1,1,14][1,-1,1,-1,-14] 2 4 $\mathbb{Z}/4\mathbb{Z}$ 0 N Y
+19a1 [0,1,1,0,9][0,1,1,0,-9] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0 Y N
+20a1 $[0,1,0,4,4]$ 2 6 $\mathbb{Z}/6\mathbb{Z}$ 0 N N
+21a1 [1,0,0,1,2][1,0,0,-1,-2] 2 8 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/4\mathbb{Z}$ 0 N Y
+24a1 [0,1,0,4,4][0,-1,0,-4,4] 2 8 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/4\mathbb{Z}$ 0 Y N
+26a1 [1,1,1,3,3][1,-1,1,-3,3] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0 N Y
+26b1 [1,1,1,56,10][1,-1,1,-56,-10] 2 7 $\mathbb{Z}/7\mathbb{Z}$ 0 N N
+27a1121212Let $h_{0}=1$ and $h_{n}$ be the coefficient of $q_{*}^{p_{n}}$ in the Fourier expansion of curve 27a1 for $n$ positive.
+Let $j_{0}=1$ and let $j_{n}$ for $n$ positive satisfy $j_{n}=nh_{n}-\sum_{r=0}^{n-1}j_{r}h_{n-r}$.
+For $n\leq 400$, we observed that $j_{n}=-2\times 2^{n/2}$ for even positive $n$, and $j_{n}=0$ for odd positive $n$.
+Obviously, we can reverse this and write down a conjecture for the Fourier expansion.
+See curve 27a1 on primes 1mar26 in [Bre25].
+(Sadly, the $j_{n}$ for $h_{n}=\tau(p_{n})$ do not appear to have similar behavior.) [0,0,1,0,7][0,0,1,0,-7] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0 N Y [^13] More precisely, it appears to us that the minima of 27a1b on primes lie on the sum of a logarithmic curve and an oscillating one, or, the sum of at least two oscillatory curves, one having a much greater wavelength than the other.
+Fourier analysis by Anthropic’s “Claude” model favors the first possibility.
+See curve 27a1 on primes 2mar26 in [Bre25].
+30a1 [1,1,0,3,2][1,-1,0,-3,2] 2 6 $\mathbb{Z}/6\mathbb{Z}$ 0
+32a1 $[0,0,0,4,0]$ 2 4 $\mathbb{Z}/4\mathbb{Z}$ 0
+33a1 [1,1,0,2,0][1,1,0,-2,0] 2 4 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 0
+34a1 [1,1,0,1,1][1,1,0,-1,-1] 2 6 $\mathbb{Z}/6\mathbb{Z}$ 0
+35a1 [0,1,1,1,2][0,1,1,1,-2] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+36a1 [0,0,0,3,2][0,0,0,-3,-2] 2 6 $\mathbb{Z}/6\mathbb{Z}$ 0
+37a1 [0,0,1,1,0][0,0,1,-1,0] 2 1 0 1 Y Y
+37b1 [0,1,1,23,50][0,1,1,-23,-50] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+38a1 [1,1,1,1,0][1,-1,1,-1,0] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+38b1 [1,1,1,4,4][1,1,1,-4,4] 2 5 $\mathbb{Z}/5\mathbb{Z}$ 0
+39a1 [1,1,0,0,2][1,-1,0,0,-2] 2 4 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 0
+40a1 [0,0,0,3,2][0,0,0,-3,2] 2 4 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 0
+42a1 [1,1,0,2,1][1,1,0,-2,-1] 2 8 $\mathbb{Z}/8\mathbb{Z}$ 0
+43a1 $[0,1,1,0,0]$ 2 1 0 1 Y Y
+44a1 [0,1,0,5,4][0,-1,0,-5,4] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+45a1 [1,1,1,1,1][1,1,1,-1,-1] 2 2 $\mathbb{Z}/2\mathbb{Z}$ 0
+46a1 [1,0,1,1,0][1,0,1,-1,0] 2 2 $\mathbb{Z}/2\mathbb{Z}$ 0
+48a1 $[0,0,0,1,0]$ 2 4 $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 0
+49a1 [1,1,0,2,0][1,-1,0,-2,0] 2 2 $\mathbb{Z}/2\mathbb{Z}$ 0
+50a1 [1,1,1,2,1][1,1,1,-2,-1] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+50b1 $[1,1,1,0,0]$ 2 5 $\mathbb{Z}/5\mathbb{Z}$ 0
+51a1 [1,1,0,0,1][1,1,0,0,-1] 2 3 $\mathbb{Z}/3\mathbb{Z}$ 0
+52a1 [0,0,0,1,2][0,0,0,1,-2] 2 2 $\mathbb{Z}/2\mathbb{Z}$ 0
+53a1 [1,1,1,8,6][1,-1,1,-8,6] 2 1 0 1 Y N
+Table 9.
+Mordell-Weil generators and Galois representation surjectivity for $\Delta$ and elliptic curves with conductor $\leq 53$.
+Column labels: A = Curve label, B = $j$-invariant, C = Generator canonical heights (approx.), D = Regulator (approx.), E = Galois rep.
+surjective for all primes?, F = Deformed plots with $h(n)=a(n)$ oscillate?
+G = Deformed plots with $h(n)=a(1+p_{n})$ oscillate?
+H = Deformed plots with $h(n)=a(p_{n})$ oscillate?
+Heights shown for rank 1 curves only.
+For rank 0 curves: regulator = 1.
+A B C D E F G H
+$\Delta$ Y N N Y
+11a1 $2^{12}$ none 1 N Y Y N
+14a1 $3^{3}\cdot 5^{3}/2^{2}$ none 1 N Y Y N
+15a1 $5^{2}\cdot 241^{3}/(2^{5}\cdot 3^{3})$ none 1 N Y Y
+17a1 $17\cdot 373^{3}/2^{17}$ none 1 N N Y
+19a1 $2^{15}/19$ none 1 N Y N
+20a1 $2^{18}/5$ none 1 N N N
+21a1 $3^{3}\cdot 5^{3}\cdot 101^{3}/(7\cdot 2^{21})$ none 1 N N Y
+24a1 $2^{12}/3$ none 1 N Y N
+26a1 $17^{3}\cdot 41^{3}/(2\cdot 13^{3})$ none 1 N N Y
+26b1 $7\cdot 11^{3}/13$ none 1 N N N
+27a1 0141414sic.
+The $j$-invariant is zero because 27a1 has complex multiplication by $\mathbb{Z}[e^{2\pi i/3}]$.
+See Theorem  III.10.1 for the classification $|\mathrm{Aut}(E)|\in\{2,4,6\}$ and its relation to $j(E)\in\{0,1728\}$ in [Si] and Theorem 5.21 (p.
+23) for the proof that $\mathrm{Aut}(E)\cong\mathbb{Z}/2\mathbb{Z}$, $\mathbb{Z}/4\mathbb{Z}$, or $\mathbb{Z}/6\mathbb{Z}$ according as $j(E)\neq 0,1728$; $j(E)=1728$; or $j(E)=0$ in [F].
+An obvious thought is that the value zero is connected to the footnotes in the row for curve 27a1 in the previous table.
+none 1 N N
+30a1 $2^{4}\cdot 5^{3}\cdot 7^{3}/(3\cdot 31)$ none 1 N
+32a1 $1728$ none 1 N
+33a1 $2^{5}\cdot 5^{3}/3$ none 1 N
+34a1 $2^{4}\cdot 17^{3}/3$ none 1 N
+35a1 $5\cdot 29^{3}/7$ none 1 N
+36a1 $2^{4}\cdot 3^{3}\cdot 5^{3}$ none 1 N
+37a1 $2^{15}\cdot 3^{3}/37$ $0.0511$ $0.0511$ Y Y Y
+37b1 $9317^{3}/37^{3}$ none 1 N
+38a1 $2^{4}\cdot 19^{3}/3$ none 1 N
+38b1 $5^{2}\cdot 127^{3}/(2\cdot 19)$ none 1 N
+39a1 $2^{4}\cdot 5^{3}\cdot 11^{3}/(3\cdot 13)$ none 1 N
+40a1 $2^{10}\cdot 5^{3}/3$ none 1 N
+42a1 $2^{4}\cdot 3^{3}\cdot 5^{3}\cdot 23^{3}/(7\cdot 43)$ none 1 N
+43a1 $2^{12}\cdot 3^{3}/43$ $0.2043$ $0.2043$ N Y Y
+44a1 $2^{4}\cdot 5^{3}/11$ none 1 N
+45a1 $5^{3}\cdot 3^{3}\cdot 2^{4}/(5\cdot 3)$ none 1 N
+46a1 $2^{4}\cdot 23^{3}/3$ none 1 N
+48a1 $2^{8}\cdot 3^{3}$ none 1 N
+49a1 $7^{5}$ none 1 N
+50a1 $2^{4}\cdot 5^{3}\cdot 7^{3}/(5\cdot 3)$ none 1 N
+50b1 $2^{4}\cdot 5^{3}\cdot 3^{3}/25$ none 1 N
+51a1 $2^{4}\cdot 5^{3}\cdot 17^{3}/(3\cdot 17)$ none 1 N
+52a1 $2^{6}\cdot 3^{3}\cdot 5^{3}/13$ none 1 N
+53a1 $2^{18}/53$ $0.2581$ $0.2581$ N Y N
+Table 10.
+Galois representation exceptional primes and modular group data for $\Delta$ and elliptic curves with conductor $\leq 53$.
+Column labels: A = Curve label, B = Galois image type (S = surjective, NS =non-surjective, CM = has complex multiplication), C = Index of $\Gamma_{0}(N)$ in $\text{SL}_{2}(\mathbb{Z})$, D = Deformed plots with $h(n)=a(n)$ oscillate?
+E = Deformed plots with $h(n)=a(1+p_{n})$ oscillate?
+F = Deformed plots with $h(n)=a(p_{n})$ oscillate?
+A B C D E F
+$\Delta$ surj. (wt 12) 1 N N Y
+11a1 NS [5] 12 Y Y N
+14a1 NS [2, 3] 24 Y Y N
+15a1 NS [2, 5] 24 Y Y
+17a1 NS [2] 24 N Y
+19a1 NS [3] 24 Y N
+20a1 NS [2, 3] 48 N N
+21a1 NS [2, 7] 48 N Y
+24a1 NS [2, 5] 48 Y N
+26a1 NS [3, 13] 48 N Y
+26b1 NS [3, 7] 48 N N
+27a1 CM 36 N
+30a1 NS [2, 3, 5] 72
+32a1 CM 48
+33a1 NS [2, 11] 72
+34a1 NS [2, 3] 72
+35a1 NS [3, 7] 72
+36a1 CM 72
+37a1 NS [37] 48 Y Y
+37b1 NS [3] 48
+38a1 NS [3, 19] 72
+38b1 NS [3, 5] 72
+39a1 NS [2, 13] 96
+40a1 NS [2, 5] 96
+42a1 NS [2, 3, 7] 144
+43a1 NS [43] 56 Y Y
+44a1 NS [2, 3] 96
+45a1 NS [2, 3] 108
+46a1 NS [2, 23] 96
+48a1 NS [2, 5] 96
+49a1 CM 84
+50a1 NS [3, 5] 120
+50b1 NS [3, 5] 120
+51a1 NS [2, 3] 96
+52a1 NS [2, 13] 144
+53a1 NS [53] 72 Y N
+Table 11.
+Modular curve genus, complex multiplication, and functional equation data for $\Delta$ and elliptic curves with conductor $\leq 53$.
+Column labels: A = Curve label, B = Genus of $X_{0}(N)$, C = Number of cusps of $X_{0}(N)$, D = Complex multiplication?, E = CM discriminant (if applicable), F = Root number of L-function, G = Deformed plots with $h(n)=a(n)$ oscillate?
+H = Deformed plots with $h(n)=a(1+p_{n})$ oscillate?
+K = Deformed plots with $h(n)=a(p_{n})$ oscillate?
+A B C D E F G H K
+$\Delta$ 0 1 N 0 +1 N N Y
+11a1 1 2 N 0 +1 Y Y N
+14a1 1 4 N 0 +1 Y Y N
+15a1 1 4 N 0 +1 Y Y
+17a1 1 4 N 0 +1 N Y
+19a1 1 3 N 0 +1 Y N
+20a1 1 6 N 0 +1 N N
+21a1 1 8 N 0 +1 N Y
+24a1 1 8 N 0 +1 Y N
+26a1 2 6 N 0 +1 N Y
+26b1 2 6 N 0 +1 N N
+27a1 1 4 Y -3 +1 N
+30a1 3 12 N 0 +1
+32a1 1 6 Y -4 +1
+33a1 3 12 N 0 +1
+34a1 3 9 N 0 +1
+35a1 3 12 N 0 +1
+36a1 3 12 Y -3 +1
+37a1 2 4 N 0 -1 Y Y
+37b1 2 4 N 0 +1
+38a1 3 9 N 0 +1
+38b1 3 9 N 0 +1
+39a1 3 16 N 0 +1
+40a1 3 12 N 0 +1
+42a1 5 24 N 0 +1
+43a1 3 4 N 0 -1 Y Y
+44a1 3 12 N 0 +1
+45a1 5 12 N 0 +1
+46a1 5 12 N 0 +1
+48a1 5 16 N 0 +1
+49a1 3 8 Y -7 +1
+50a1 5 15 N 0 +1
+50b1 5 15 N 0 +1
+51a1 5 12 N 0 +1
+52a1 5 18 N 0 +1
+53a1 5 6 N 0 -1 Y N
+Figure Jupyter SageMath notebook
+Figure 1 tauprime=h logarithmic envelopes prec100 n400 1dec25
+Figure 2 tauprime c=2 4feb26
+Figure 3 tauprime c=2 4feb26
+Figure 4 tauprime c=3 4feb26
+Figure 5 tauprime many c’s 18feb26
+Figure 6 tauprime controls 22feb26
+Figure 7 undeformed primeTau 5oct25
+Figure 8A curve 11a1 29jan26
+Figure 8B curve 11a1 $h(n)=a(p_{n})$ 9mar26
+Figure 8C curve 11a1 $h(n)=a(p_{n}+1)$ 8mar26
+Figure 9A curve 14a1 9feb26
+Figure 9B curve 14a1 $h(n)=a(p_{n})$. 9mar26
+Figure 9C curve 14a1 $h(n)=a(p_{n}+1)$. 9mar26
+Figure 10A curve 15a1 10feb26
+Figure 10B curve 15a1 on primes 10feb26
+Figure 11A curve 17a1 30jan26
+Figure 11B curve 17a1 on primes 2feb26
+Figure 12A curve 19a1 31jan26
+Figure 12B curve 19a1 on primes 21feb26
+Figure 13A curve 20a1 11feb26
+Figure 13B curve 20a1 on primes 10feb26
+Figure 14A curve 21a1 11feb26
+Figure 14B curve 21a1 on primes 27feb26
+Figure 15A curve_24a1 all indices 21feb26
+Figure 15B curve 24a1 on primes 23feb26
+Figure 16A curve 26a1 28feb26
+Figure 16B curve 26a1 on primes 28feb26
+Figure 17A curve 26b1 28feb26
+Figure 17B curve 26b1 on primes
+Figure 18A crv 27a1 28feb26
+Figure 18B curve 27a1 on primes 1mar26no2
+Figure 19A curve 37a1 1feb26
+Figure 19B curve 37a1 on primes 23feb26
+Figure 20A curve 43a1no2 22feb26
+Figure 20B curve 43a1 on primes 28feb26
+Figure 21A crv 53a1 2feb26
+Figure 21B curve 53a1 on primes 28feb26
+Figure 22 deformed tau 18oct25
+Table 12. Figure references. The notebooks are in our repository [Bre25].
+8. Figures
+Figure 1. Minimum moduli for $h(n)=\tau(p_{n})$ with $c=1$.
+Figure 2. Minimum moduli and their logarithms for $h(n)=\tau(p_{n})$ with $c=2$.
+Figure 3. Before and after de-trending for $h(n)=\tau(p_{n})$ with $c=2$.
+Figure 4. Minimum moduli and their logarithms for $h(n)=\tau(p_{n})$ with $c=3$.
+Figure 5. Minimum moduli over many values of $c$ for $h(n)=\tau_{p}(n)$.
+Figure 6. Experimental control: minimum moduli for $h(n)=\tau(p_{n}+1)$ with $c=1$.
+Figure 7. Minimum moduli for $h(n)=\tau(p_{n})$ before deformation.
+(a) $h(n)=a_{n}$.
+(b) $h(n)=a_{p_{n}}$.
+(c) $h(n)=a_{1+p_{n}}$.
+Figure 8. Curve 11a1 with $c=1$.
+(a) $h(n)=a_{n}$.
+(b) $h(n)=a_{p_{n}}$.
+(c) $h(n)=a_{1+p_{n}}$.
+Figure 9. Curve 14a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 10. Curve 15a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 11. Curve 17a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 12. Curve 19a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 13. Curve 20a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 14. Curve 21a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 15. Curve 24a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 16. Curve 26a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 17. Curve 26b1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 18. Curve 27a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 19. Curve 37a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 20. Curve 43a1 with $c=1$.
+(a) $h(n)=a(n)$.
+(b) $h(n)=a_{1+p_{n}}$.
+Figure 21. Curve 53a1 with $c=1$.
+Figure 22. Minimum moduli for $h(n)=\tau(n)$ with $c=0$.